OpenStudy (anonymous):

help with 2 step equations?

OpenStudy (anonymous):

ex.?

OpenStudy (anonymous):

50=2x+21

OpenStudy (anonymous):

you are trying to isolate x. so first move the 21 to the other side by subtracting 21 from each side: 50-21=2x+21-21

OpenStudy (anonymous):

simplify that part out.

OpenStudy (anonymous):

x=29/2

OpenStudy (anonymous):

not skipping steps rob. showing how to work the problem not just give the answer. 50-21=29, 2x+21-21=2x 29=2x divide both sides by 2 to isolate the x 29/2=x